# Quantum Enhanced Sensitivity through Many-Body Bloch Oscillations

### v1

@@ Abstract
Effect of excitations.—In the previous section, we focused on the half-filling case, namely $N{=}L/2$. In this section, we investigate the impact of the number of excitations in more details. For the sake of symmetry, we consider odd system sizes $L$ and choose the initial states to be symmetric around the center of the chain. For instance, if the system size is $L{=}7$ then the initial state with $N=2$ and $N=3$ excitations are set to $\mathinner{|{0,0,1,0,1,0,0}\rangle}$ and $\mathinner{|{0,1,0,1,0,1,0}\rangle}$, respectively. In Figs. 4(a) and 4(b) the normalized QFI is plotted for different excitation numbers $N$ in a system of size $L{=}17$ for $\Delta=0$ and $\Delta=1$, respectively. Notably, increasing the excitations always enhances the QFI, indicating better sensitivity. In addition, in the localized phase the QFI shows excitation-dependence. In order to discriminate the QFI enhancement due to the system size from the excitation number, in Figs. 4(c) and 4(d) we plot the long-time behavior of $\mathcal{F}_{Q}/t^{2}$ as a function of $h/J$ for a fixed number of excitation $N{=}3$ for various choices of $L$. Interestingly, the size dependence completely disappears for $h>h_{c}$ indicating that the sensitivity is only enhanced through the number of excitations when the system is in the localized phase (The scaling of the normalized QFI at the transition points in Fig. 4(c) and (d) are discussed in SM). These observations suggest the following ansatzes for the QFI in many-body probes with multiple excitations
$\mathcal{F}_{Q}$ $\sim h^{-2}t^{2}L^{0}N^{\alpha},\;\;\text{ for }h>h_{c}~{},$ (4)
$\mathcal{F}_{Q}$ $\sim t^{2}L^{\beta}N^{\alpha},\;\;\;\;\;\;\;\;\text{ for }h\leq h_{c}~{},$ (5)
@@
Conclusion.—We have studied the metrological aspects of the Bloch oscillations in single- and many-particle Stark systems which support a phase transition, across the entire spectrum, from an extended to a localized phase. Unlike conventional equilibrium criticality-based quantum sensors, our non-equilibrium probe does not require complex initialization and achieves quantum enhanced sensitivity across the entire extended phase. In addition, our analysis provides an ansatz for the QFI in terms of time, probe size, and number of excitations. In the long-time limit, the QFI scales quadratically with time. In contrast, the precision with respect to the probe size transforms from super-Heisenberg scaling in the extended phase to size-independent behavior in the localized phase. Finally, the number of excitations always enhances the precision of the probe, though, the amount of enhancement depends on the interaction between excitations. Given its time-dependent feature, our work sets the groundstone for further precision measurement experiments, e.g. in gravitometery or electric field sensing.
Acknowledgements.
@@
Acknowledgements.—The authors thank the Center for High Performance Computing at Shanghai Jiao Tong University for providing access to the Siyuan-1 cluster. The work of MA was partially supported by the Yangyang Development Foundation. AB acknowledges support from the National Natural Science Foundation of China (Grants No. 12050410253, No. 92065115, and No. 12274059), and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). HM acknowledges partial financial support by Iran Optics and Quantum Technologies Development Council.
Supplemental Material:
Quantum enhance sensitivity through many-body Bloch oscillations
In this supplemental material we provide further information and more details on the variations of the quantum Fisher information (QFI) both in single- and multi-particle cases. We also present additional information on the Bloch oscillations in the many-body case.
S1 Single particle QFI scaling
@@
Further details in studying the size-dependence of the QFI that have brought us to Eqs. (2) in the main text are presented in Fig. S1 where we plot the long-time normalized QFI as a function of system size $L$ for different choices of $h$. Indeed, for any given $h$, one observes a power law scaling growth of $\mathcal{F}_{Q}/t^{2}$ with respect to $L$ which experiences an enhancement before turning into a plateau around the size $L{\simeq}8J/h_{c}$. Since the transition point depends on the system size the plateau sets on at different $L$ values. One can clearly identify an improved scaling with size close to the transition point.
@@
Figure S1: The normalized QFI for different gradient fields $h$ is plotted as a function of system size $L$. The plateau for a specific $h$ resembles the localized phase in Fig. 2(a) in the main text.
S2 Many-Body Bloch oscillation
@@
In this section, we present the results of the Bloch oscillations in the many-body case to compare with the single excitation in Fig. 1 in the main text. To see the behavior of the many-body excitation oscillations at different gradient fields, we compute the dynamics of the probability of having an excitation in every site of the system $P_{l}(t)$. The results are presented in Fig. S2 for the system size $L=11$ with the initial state $\mathinner{|{\Psi_{0}}\rangle}=\mathinner{|{01010101010}\rangle}$. Figs. S2(a)-(c) are plotted for the non-interacting sites ($\Delta=0$) in the extended phase, at the transition point and in the localized phase, respectively. In particular, in Figs. S2(a) the extended phase at $h/J=0.05$ shows only partial excitation revivals for a few sites at the middle of the lattice. However, as in Fig. S2(b) which shows the probabilities at the transition point $h_{c}/J=0.4$ (extracted from Fig. 3(a) in the main text) complete revival for all sites are visible at particular times, e.g. at $tJ=65$. Meanwhile, Fig. S2(c) corresponds to $h/J=5$ which represents a system in the localized phase.
@@
In Figs. S2(d)-(f) the same analysis performed for the homogeneous Heisenberg model ($\Delta=1$) are presented for the same initial state. Even though the localized phase exhibits a behavior akin to the non-interacting case, dynamics of the probabilities in the extended phase and at the transition points are dramatically different. As shown in Fig. S2(e) in the transition point where $h_{c}/J=0.9$ (see Fig. 3(b) in the main text), a clear separation between ground- and excited-states is obvious. Comparing Figs. S2(b) and (e) at the transition points, we can deduce that by increasing the interaction strength $\Delta$, the particle excitations tend to occupy sites with higher energies, emerging a clear separation between the excited sites and those who remain in the ground state, see Fig. S2(e).
@@
Figure S2: The probability distribution of each site $P_{l}$ for the initial state $\mathinner{|{\Psi_{0}}\rangle}=\mathinner{|{01010101010}\rangle}$ are depicted for the system size $L=11$ for $\Delta=0$ at (a) $h/J=0.05$ , (b) $h/J=0.4$ and (c) $h/J=5$, in extended phase, at the transition point, and in the localized phase, respectively. In (d)-(f) $P_{l}$ are shown for the Heisenberg Hamiltonian $\Delta=1$ at $h/J=0.05$ in the extended phase, $h/J=0.9$ at the transition point, and $h/J=5$ in the localized phase, respectively
S3 Many-body QFI scaling
In studying the single-particle case, we have shown in the main text that the QFI scales as $\mathcal{F}_{Q}\sim t^{2}L^{\beta}$ where $\beta$ depends on the phase of the system, which in turn, is determined by $h$ and $L$ [see Fig. 2(c) in the main text]. In particular, we find $\beta{=}2$ at the transition point. Meanwhile, in the many-body case depending on the number of excitations in the initial state we find the scaling function as $\mathcal{F}_{Q}\sim t^{2}L^{\beta}N^{\alpha}$. Unlike $\beta$ which dramatically changes with the phase, $\alpha$ is almost independent of $h$. This behavior especially is observed near the transition point. Therefore, for a fixed number of excitations the QFI scales as $\mathcal{F}_{Q}(h_{c})\sim t^{2}L^{\beta}$ with $\beta\approx 2$. To clearly show this point, in Figs. S3(a) and (b) for both $\Delta=0$ and $\Delta=1$ we plot the normalized QFI at the transition points $h=h_{c}$ for a fixed number of excitations ($N{=}3$) and different system sizes $L$. The fitting lines confirm the power law scaling $\mathcal{F}_{Q}(h_{c})/t^{2}\sim L^{\beta}$ with $\beta=2.1$ for $\Delta=0$ and $\beta=2.2$ for $\Delta=1$, which are very close to the value obtained from single-particle case $\beta=2$.
@@
Figure S3: the normalized QFI at the transition point $\mathcal{F}_{Q}(h_{c})/t^{2}$ for excitation number $N=3$ as a function of $L$ for (a) $\Delta=0$ and (b) $\Delta=1$. The best fitting exponents for $\mathcal{F}_{Q}(h_{c})/t^{2}\sim L^{\beta}$ are $\beta=2.1$ for $\Delta=0$ and $\beta=2.2$ for $\Delta=1$.

### v2

@@ 4 Model 2: many-body probe (half filling)
5 Effect of excitations
In the previous section, we focused on the half-filling case, namely $N{=}L/2$. In this section, we investigate the impact of the number of excitations in more details. For the sake of symmetry, we consider odd system sizes $L$ and choose the initial states to be symmetric around the center of the chain. For instance, if the system size is $L{=}7$ then the initial state with $N=2$ and $N=3$ excitations are set to $\mathinner{|{0,0,1,0,1,0,0}\rangle}$ and $\mathinner{|{0,1,0,1,0,1,0}\rangle}$, respectively. In Figs. 4(a) and 4(b) the normalized QFI is plotted for different excitation numbers $N$ in a system of size $L{=}17$ for $\Delta=0$ and $\Delta=1$, respectively. Notably, increasing the excitations always enhances the QFI, indicating better sensitivity. In addition, in the localized phase the QFI shows excitation-dependence. In order to discriminate the QFI enhancement due to the system size from the excitation number, in Figs. 4(c) and 4(d) we plot the long-time behavior of $\mathcal{F}_{Q}/t^{2}$ as a function of $h/J$ for a fixed number of excitation $N{=}3$ for various choices of $L$. Interestingly, the size dependence completely disappears for $h>h_{c}$ indicating that the sensitivity is only enhanced through the number of excitations when the system is in the localized phase (The scaling of the normalized QFI at the transition points in Fig. 4(c) and (d) are discussed in Appendix C). These observations suggest the following ansatzes for the QFI in many-body probes with multiple excitations
$\mathcal{F}_{Q}$ $\sim h^{-2}t^{2}L^{0}N^{\alpha},\;\;\text{ for }h>h_{c}\leavevmode\nobreak\ ,$ (7)
$\mathcal{F}_{Q}$ $\sim t^{2}L^{\beta}N^{\alpha},\;\;\;\;\;\;\;\;\text{ for }h\leq h_{c}\leavevmode\nobreak\ ,$ (8)
@@ 5 Effect of excitations
6 Practical implementation
As discussed before, QFI quantifies the ultimate precision limit in a sensing procedure which can only be achieved by an optimal measurement. In practice, however, due to practical constraints performing such an optimal measurement might be very challenging. In this section, we consider simple available measurements in both single- and multi-particle cases. For the single excitation sensors, we consider a simple position measurement described by projectors $\{\pi_{l}{=}\mathinner{|{l}\rangle}\mathinner{\langle{l}|}\}$, where $l{=}1,2,\cdots,L$. Therefore, the probability of each outcome is $p_{l}(t){=}|\braket{l}{\Psi(t)}|^{2}$. The obtainable precision by this measurement is bounded by the CFI, see Eq. (1). In Fig. 6(a), we compare the long-time behavior of the CFI and the QFI in a single-excitation probe of size $L=16$ as a function of $h/J$. As the figure shows, the CFI qualitatively follows the QFI, though it always takes lower values. In the inset, we plot the ratio $\mathcal{F}_{C}/\mathcal{F}_{Q}$ as a function of $h/J$ which shows that our simple measurement is almost optimal in the extended phase (i.e. $\mathcal{F}_{C}/\mathcal{F}_{Q}\simeq 1$) and it becomes sub-optimal (i.e. $\mathcal{F}_{C}/\mathcal{F}_{Q}\sim 0.5$) as localization strength is enhanced. Interestingly, the lowest ratio of $\mathcal{F}_{C}/\mathcal{F}_{Q}$ takes place around the transition point showing that our simple measurement scheme cannot efficiently extract information from the complex critical quantum state.
In the case of a multi-particle probe, we measure the configuration of particles which is obtained by a simple particle detection imaging. The measurement can be described by $\{\pi_{{}_{\mathbf{Z}}}{=}\mathinner{|{\mathbf{z}}\rangle}\mathinner{\langle{\mathbf{z}}|}\}$, where $\mathinner{|{\mathbf{z}}\rangle}$ accounts for all possible particle configurations in the probe. For instance, in a probe of size $L$ with $N$ particles (i.e. excitations), there are (${L\choose N}$ different configurations. In the case of $N{=}1$, the problem reduces to the position measurement, discussed above. Each configuration is observed by the probability $p_{{}_{\mathbf{z}}}(t){=}|\braket{\mathbf{z}}{\Psi(t)}|^{2}$, for which one can obtain the corresponding CFI. In Fig. 6(b), we compare the long-time behavior of the CFI and the QFI in a half-filled probe of size $L=16$ (with $N=L/2$ excitations) initialized in the Neel state as a function of $h/J$. Similar to the case of single excitation, the CFI qualitatively follows the QFI, though with lower values. In the inset, we plot $\mathcal{F}_{C}/\mathcal{F}_{Q}$ as a function of $h/J$. Similar to the single excitation case, it shows that a simple configuration measurement is almost optimal in the extended phase and becomes sub-optimal with $\mathcal{F}_{C}/\mathcal{F}_{Q}\simeq 0.5$ as $h/J$ increases. Again the lowest ratio of $\mathcal{F}_{C}/\mathcal{F}_{Q}$ is obtained around the transition point where our simple measurement scheme cannot fully extract the information from the complex critical quantum state of the system.
(a)Refer to caption(b)
Figure 6: The long-time limit of $\mathcal{F}_{C}$ (purple solid lines) and $\mathcal{F}_{Q}$ (green dashed lines) as a function of $h/J$ for the system size $L=16$ are plotted for the initial state of: (a) a single excitation state $\mathinner{|{1}\rangle}$; and (b) the half-filled Neel state with $\Delta=0$. The insets, depicting the ratio of $\mathcal{F}_{C}/\mathcal{F}_{Q}$ versus $h/J$, show that our specific choice of simple measurements can reach around $50\%$ of the ultimate precision bound, determined by the QFI.
7 Conclusion
@@
We have studied the metrological aspects of the Bloch oscillations in single- and many-particle Stark systems which support a phase transition, across the entire spectrum, from an extended to a localized phase. Unlike conventional equilibrium criticality-based quantum sensors, our non-equilibrium probe does not require complex initialization and achieves quantum enhanced sensitivity across the entire extended phase. In addition, our analysis provides an ansatz for the QFI in terms of time, probe size, and number of excitations. In the long-time limit, the QFI scales quadratically with time in both extended and localized phases. In contrast, the precision with respect to the probe size transforms from super-Heisenberg scaling in the extended phase to size-independent behavior in the localized phase. Finally, the number of excitations always enhances the precision of the probe, though, the amount of enhancement depends on the interaction between excitations. We also show that a simple configuration measurement can achieve an estimation accuracy which is not far from the ultimate precision limit. Given its time-dependent feature, our work sets the groundstone for further precision measurement experiments, e.g. in gravitometery or electric field sensing.
Acknowledgements
@@
The authors thank the Center for High Performance Computing at Shanghai Jiao Tong University for providing access to the Siyuan-1 cluster. The work of MA was partially supported by the Yangyang Development Foundation. AB acknowledges support from the National Natural Science Foundation of China (Grants No. 12050410253, No. 92065115, and No. 12274059), and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). HM acknowledges partial financial support by Iran Optics and Quantum Technologies Development Council.
Appendix A Appendix A: Single particle QFI scaling
@@
Further details in studying the size-dependence of the QFI that have brought us to Eqs. (5) in the main text are presented in Fig. 7 where we plot the long-time normalized QFI as a function of system size $L$ for different choices of $h$. Indeed, for any given $h$, one observes a power law scaling growth of $\mathcal{F}_{Q}/t^{2}$ with respect to $L$ which experiences an enhancement before turning into a plateau around the size $L{\simeq}8J/h_{c}$. Since the transition point depends on the system size the plateau sets on at different $L$ values. One can clearly identify an improved scaling with size close to the transition point.
@@
Figure 7: The normalized QFI for different gradient fields $h$ is plotted as a function of system size $L$. The plateau for a specific $h$ resembles the localized phase in Fig. 2(a) in the main text.
Appendix B Appendix B: Many-Body Bloch oscillation
@@
In this section, we present the results of the Bloch oscillations in the many-body case to compare with the single excitation in Fig. 1 in the main text. To see the behavior of the many-body excitation oscillations at different gradient fields, we compute the dynamics of the probability of having an excitation in every site of the system $P_{l}(t)$. The results are presented in Fig. 8 for the system size $L=11$ with the initial state $\mathinner{|{\Psi_{0}}\rangle}=\mathinner{|{01010101010}\rangle}$. Figs. 8(a)-(c) are plotted for the non-interacting sites ($\Delta=0$) in the extended phase, at the transition point and in the localized phase, respectively. In particular, in Figs. 8(a) the extended phase at $h/J=0.05$ shows only partial excitation revivals for a few sites at the middle of the lattice. However, as in Fig. 8(b) which shows the probabilities at the transition point $h_{c}/J=0.4$ (extracted from Fig. 3(a) in the main text) complete revival for all sites are visible at particular times, e.g. at $tJ=65$. Meanwhile, Fig. 8(c) corresponds to $h/J=5$ which represents a system in the localized phase.
@@
In Figs. 8(d)-(f) the same analysis performed for the homogeneous Heisenberg model ($\Delta=1$) is presented for the same initial state. Even though the localized phase exhibits a behavior akin to the non-interacting case, dynamics of the probabilities in the extended phase and at the transition points are dramatically different. As shown in Fig. 8(e) in the transition point where $h_{c}/J=0.9$ (see Fig. 3(b) in the main text), a clear separation between ground- and excited-states is obvious. Comparing Figs. 8(b) and (e) at the transition points, we can deduce that by increasing the interaction strength $\Delta$, the particle excitations tend to occupy sites with higher energies, emerging a clear separation between the excited sites and those who remain in the ground state, see Fig. 8(e).
@@
Figure 8: The probability distribution of each site $P_{l}$ for the initial state $\mathinner{|{\Psi_{0}}\rangle}=\mathinner{|{01010101010}\rangle}$ are depicted for the system size $L=11$ for $\Delta=0$ at (a) $h/J=0.05$ , (b) $h/J=0.4$ and (c) $h/J=5$, in extended phase, at the transition point, and in the localized phase, respectively. In (d)-(f) $P_{l}$ are shown for the Heisenberg Hamiltonian $\Delta=1$ at $h/J=0.05$ in the extended phase, $h/J=0.9$ at the transition point, and $h/J=5$ in the localized phase, respectively
Appendix C Appendix C: Many-body QFI scaling
In studying the single-particle case, we have shown in the main text that the QFI scales as $\mathcal{F}_{Q}\sim t^{2}L^{\beta}$ where $\beta$ depends on the phase of the system, which in turn, is determined by $h$ and $L$ [see Fig. 2(c) in the main text]. In particular, we find $\beta{=}2$ at the transition point. Meanwhile, in the many-body case depending on the number of excitations in the initial state we find the scaling function as $\mathcal{F}_{Q}\sim t^{2}L^{\beta}N^{\alpha}$. Unlike $\beta$ which dramatically changes with the phase, $\alpha$ is almost independent of $h$. This behavior especially is observed near the transition point. Therefore, for a fixed number of excitations the QFI scales as $\mathcal{F}_{Q}(h_{c})\sim t^{2}L^{\beta}$ with $\beta\approx 2$. To clearly show this point, in Figs. 9(a) and (b) for both $\Delta=0$ and $\Delta=1$ we plot the normalized QFI at the transition points $h=h_{c}$ for a fixed number of excitations ($N{=}3$) and different system sizes $L$. The fitting lines confirm the power law scaling $\mathcal{F}_{Q}(h_{c})/t^{2}\sim L^{\beta}$ with $\beta=2.1$ for $\Delta=0$ and $\beta=2.2$ for $\Delta=1$, which are very close to the value obtained from single-particle case $\beta=2$.
@@ Appendix C Appendix C: Many-body QFI scaling
Figure 9: the normalized QFI at the transition point $\mathcal{F}_{Q}(h_{c})/t^{2}$ for excitation number $N=3$ as a function of $L$ for (a) $\Delta=0$ and (b) $\Delta=1$. The best fitting exponents for $\mathcal{F}_{Q}(h_{c})/t^{2}\sim L^{\beta}$ are $\beta=2.1$ for $\Delta=0$ and $\beta=2.2$ for $\Delta=1$.